A tree of height 4.5 m casts a shadow of 3.0 m then the height ,in metres, of a tree casting a 2.8 m shadow is
Answer by nerdybill(7384)

You are dealing with proportions.
.
Let x = height of tree (in meters)
3/4.5 = 2.8/x
Cross multiplying we get:
3x = (4.5)(2.8)
3x = 12.6
x = 12.6/3
x = 4.2 meters